OpenStudy (anonymous):

f(x)= sqrt of 2x -8 whats x

OpenStudy (anonymous):

incomplete

Parth (parthkohli):

You can solve for \(x\).

Parth (parthkohli):

For example,\[f(x) = \sqrt{x - 8} \]Let \(f(x) = y\).\[ y = \sqrt{x - 8}\]\[y^2 = x - 8 \]\[ y^2 + 8 = x\]

OpenStudy (anonymous):

x is a variable?